OpenStudy (anonymous):

9th grade algebra help!

OpenStudy (some.random.cool.kid):

i can try

OpenStudy (anonymous):

3. The table below represents the displacement of a duck from its nest as a function of time. Time (x) Displacement (y) (hour) from nest ( Miles) 0 7 1 29 2 51 3 73 4 95 Part A: What is the y-intercept of the function and what does this tell you about the duck? (4 points) Part B: Calculate the average rate of change of the function represented by the table between x=2 to x=4 hours? (4 points) Part C: What would be the domain of the function if the duck continued to walk at this rate until it traveled 227 miles from the nest? ( 2 points) Part A: What is the y-intercept? What does this tell you about the duck? Part B: What is the average rate of change from x = 2 to x = 4? Show your work. Part C: Write a compound inequality of the domain if he traveled 227 miles from the nest?

OpenStudy (some.random.cool.kid):

forget it your on your own buddy :b nah just kidding give me a second

OpenStudy (anonymous):

lol you scared me there a second haha

OpenStudy (some.random.cool.kid):

did they give you a graph to work with or no because this looks like something you need a graph for..

OpenStudy (anonymous):

no that's all it

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

well your y intercept looks like the Displacement, correct?

OpenStudy (anonymous):

yes

OpenStudy (some.random.cool.kid):

ok so then your y intercept will be whats under the Displacement

OpenStudy (anonymous):

ok

OpenStudy (some.random.cool.kid):

so what do you think these numbers infer to you about the duck?

OpenStudy (anonymous):

I don't know

OpenStudy (some.random.cool.kid):

perhaps it tell you that he travels far? -note- im not giving away answer just trying to see if you can figure this out

OpenStudy (anonymous):

I've been stuck on this for almost a month, im trying my hardest lol

OpenStudy (some.random.cool.kid):

you waited a month to ask for help and isnt this due? :O

OpenStudy (anonymous):

no she said before winter break we have to turn it in

OpenStudy (some.random.cool.kid):

oh ok brb i come back in like 3 minutes

OpenStudy (anonymous):

ok

OpenStudy (some.random.cool.kid):

@Here_to_Help15

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

@Here_to_Help15 help her real quick i gtg got to do a test

OpenStudy (here_to_help15):

y-intercept is the y-value when x = 0. From the table, what is the y-value when x = 0?

OpenStudy (anonymous):

what?

OpenStudy (here_to_help15):

What is the answer?

OpenStudy (here_to_help15):

It is 7

OpenStudy (here_to_help15):

Ok? you got that

OpenStudy (anonymous):

yeah

OpenStudy (here_to_help15):

y-intercept = 7 What does this tell you about the duck? At time x = 0, the duck is 7 units away from the nest. That means the duck was NOT in its nest at time x = 0.

OpenStudy (anonymous):

oh

OpenStudy (here_to_help15):

Part B: What is the average rate of change from x = 2 to x = 4? Show your work. Average rate of change = y^2−y^1/x^2−x^1 Read the y-values from the table when x = 2 and when x = 4 Average rate of change = 95−51/4−2=?

OpenStudy (here_to_help15):

dam thing wont let you put terms

OpenStudy (anonymous):

hey I have to go anyways thankas a lot to both of you!

OpenStudy (here_to_help15):

Part B: What is the average rate of change from x = 2 to x = 4? Show your work. Average rate of change = y^2 y^1/x^2 x^1 Read the y-values from the table when x = 2 and when x = 4 Average rate of change = 51/4

OpenStudy (here_to_help15):

95-51/ 4-2

OpenStudy (here_to_help15):

i meant

OpenStudy (here_to_help15):

=?

OpenStudy (here_to_help15):

44 / 2 = 22.

OpenStudy (anonymous):

ok

OpenStudy (here_to_help15):

C) Slope = 22 (same as average rate of change found in part B.) y = mx + b y = 22x + b when x = 0, y = 0 + b y = b From the table, when x = 0, y = 7. Therefore, b = 7 y = 22x + 7 Part C: Write a compound inequality of the domain if he traveled 227 miles from the nest? Set y = 227 in y = 22x + 7 227 = 22x + 7 227 - 7 = 22x 22x = 220 x = 10 Domain is x = 0 to x = 10. Or, domain is: [0, 10].

OpenStudy (anonymous):

lol you type fast!

OpenStudy (here_to_help15):

yea trying to hurry before you leave

OpenStudy (anonymous):

lol thank you alot

OpenStudy (here_to_help15):

Compound inequality for the domain: \[0 \le x \le 10\]